OpenStudy (anonymous):

Use the slope formula to find the slope of the line passing through the given points. (–4, 7) and (0, 8)

rebeccaxhawaii (rebeccaxhawaii):

okay so it will be

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ y2-y1 }{ x2-x1 }\]

rebeccaxhawaii (rebeccaxhawaii):

thats your fomula

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ 8-7 }{ 0+4 }\]

rebeccaxhawaii (rebeccaxhawaii):

so then that will be 1/4

rebeccaxhawaii (rebeccaxhawaii):

the rate of change or slope is 1/4

OpenStudy (anonymous):

can you help with one more

rebeccaxhawaii (rebeccaxhawaii):

sure

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Explain why the triangles below are proportional to each other.

rebeccaxhawaii (rebeccaxhawaii):

the bigger triangle is double the size of the smaller

OpenStudy (anonymous):

Do you think you could do this. What is the slope of this line? How much does each pencil cost?

OpenStudy (anonymous):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

okay well give me two points from that graph

rebeccaxhawaii (rebeccaxhawaii):

anyone of them

OpenStudy (anonymous):

(1,0.15) and (2,0.30)

rebeccaxhawaii (rebeccaxhawaii):

okay so

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ 0.30-0.15 }{ 2-1 }\]

rebeccaxhawaii (rebeccaxhawaii):

so \[\frac{ .15 }{ 1 } = 0.15\]

rebeccaxhawaii (rebeccaxhawaii):

therefore youre slope is .15

OpenStudy (anonymous):

can you help me with all of them :)

rebeccaxhawaii (rebeccaxhawaii):

open a new question and tag me

OpenStudy (anonymous):

k